IMACS progress report no. 60
B. Bigelow 9 April 2003
I) Project Status
Currently scheduled design/fabrication tasks: Status:
a) DSS drive integration (Birk) Done
b) DSS testing (Birk/Hare) IP
c) Calibration sys. fab (Kowal) IP
d) CCD array testing (Thompson/Burley) IP
e) Dewar/Detector final integration (Thompson/Burley) IP
f) Grism procurement (Sutin) IP
g) Instrument electronics assembly (Bagish) on sched.
h) Control software design (Birk) on sched.
i) Short Camera opto-assembly (Bigelow) Re-work
j) IFU design and build (U. Durham) IP
Project events in March:
- First light with short camera!
- Thermal testing delayed until final assembly complete.
- Thermal testing enclosure test assembly completed.
- IMACS pre-ship review delayed, currently being rescheduled for late May.
- Magellan engineering time set aside for IMACS install - August 6-18, 2003.
- Magellan FC/ACD optics delayed until at least September, '03.
Latest project images:
Final Assy: http://www.ociw.edu/instrumentation/imacs/devel/final/index.html
Disperser Server: http://www.ociw.edu/instrumentation/imacs/devel/dss/index.html
Filter Servers: http://www.ociw.edu/instrumentation/imacs/devel/fss/index.html
II) Optics (Sutin)
- Echellette prism due from Zygo 5/7/03
- Monitoring progress on 3rd fabrication effort for 300l/mm grism
- Investigating ThAr calibration lamps
- Investigating project documentation programs - Docbook?
III) Electronics (Bagish)
- Took cable wrap off and re-laid cables into two wraps, re-installed wraps.
- 10 structure RTD thermal sensors attached and operating.
- Black ABS box complete and mounted. Missing one audio power connector and
test of RS-422 communications.
- Science and guider camera power supplies installed, hooked up and science
camera supplies tested through wrap.
- Pressure transducer and RS-422 communications chassis completed and
installed in rack 2.
- Ordered last (?!) set of connectors to change out shutter motors to same type
connectors (for ease of trouble-shooting).
- Attach remaining RTD thermal sensors and pressure transducers.
- Rearrange inputs to data acquisition unit.
- Mount 2nd data acquisition unit in rack 2.
- Test RS-422 communications.
- Complete calibration lamps cable.
IV) Software (Birk/Clardy)
- 4 of 7 DSS positions now mapped into control s/w
- Mapping of f/2 and f/4 camera filter servers in progress
- In Chile 4/6 - 4/17.
- Making progress on astronomer GUI for mask cutting software
- Continued de-bugging of mask cutting software with LCO staff
- Investigating guiding control issues posed by atmospheric dispersion
V) Detector Systems (Thomson/Burley)
The revisions to the PCI + DSP pcbs are almost complete, there is a
bit of clean-up and double-checking to do.
On the CCD array, we looked at a problem with charge trailing in the
horizontal direction for severely saturated images -- adjusting one of
the clock voltages seemed to fix this.
There is also some problem with the image shapes - saturated images
show wings in the horizontal direction, maybe non-saturated images do
as well. We are looking into this as it looks to be a ccd or
VI) Mechanics (Bigelow/Hare)
1) Continued plumbing installation
2) Installed second IGUS chain, re-route plumbing and cabling
3) Installed first grism, in cell, on DSS wheel
4) Continued final spacing measurements on collimator focus, cameras, dewar, BFD
5) Designed shutter motor guards, drafting done, waiting for material
6) Continue flexure testing
7) Started Re-mount of S04 in new cell
1) Re-design SMS mounting plate for new CFG layout
2) Re-assembly of short camera with new S04 cell
3) Re-test hydro-static flexure in Collimator doublet
4) Design instrument air over-pressure system (dust protection)
5) Complete design of air supply and exhuast plumbing
6) Explore new designs for slit mask frames for added stiffness
- Replaced damaged linear motor for f/2 camera shutter
- DSS wheel and drive testing under full load continuing
- New CFG design complete, drafting in progress, solids to Vince
- Minor re-design of SHG linear guides in progress
- New limit switches and mounting brackets in design for all guider stages
- Assembly of calibration lamp unit in progress
VII) Fabrication (Kowal)
1) Six grating cells completed.
2) G.T.M. Base Plate was repaired.
3) Delrin filter server guides and spacers are done.
4) Parts for second imaging mirror support are finished.
5) Calibration lamp reflector block is done.
6) Cable Tray Assembly is 75% finished.
7) Short camera barrel spacer is in progress.
Expectations for April:
1) Finish all Cable Tray parts.
2) Make new Center Field Guider parts.
3) Rework and modifications of parts as needed.
IX) Other Business
Next meeting: 10:15 AM, 5/7/03